OpenStudy (anonymous):

y + 2x = −1 3y − x = 4

OpenStudy (anonymous):

elimination: y+2x=-1 6y-2x=8 7y=7 y=1 1+2x=-1 2x=-2 x=-1 Check 3(1)-(-1)=4 3+1=4